OpenStudy (deadshot):

Solve for x: 2 over 5 plus 3 over 5 x equals quantity x plus 5 over 10?

OpenStudy (anonymous):

\[\frac{2}{5}+\frac{3}{5x}=\frac{x+5}{10}\]??

OpenStudy (deadshot):

yes

OpenStudy (deadshot):

I've tried solving it, but I don't think there is a solution

OpenStudy (deadshot):

@campbell_st can you help me with this?

OpenStudy (campbell_st):

I'd multiply every term by 10x the common denominator of the 3 fractions \[\frac{2}{5} \times 10x+ \frac{3}{5x} \times 10x = \frac{x +5}{10} \times 10x\] this can be simplified to \[4x + 6 = x(x + 5)\] which is the same as \[4x + 6 = x^2 + 5x\] or \[x^2 + x - 6 = 0\] this is a quadratic that can be solved by factoring hope this helps

OpenStudy (deadshot):

Thanks! The answer is (x-3)(x+2), right?

OpenStudy (campbell_st):

not quite you have (x -3)(x + 2) = 0 so what values can x take..?

OpenStudy (deadshot):

x can take -3 and 2, right?

OpenStudy (campbell_st):

nope... reverse the signs... because if x = 3 then (3 - 3) = 0 and 0 * (3 + 2) = 0 same for x = -2 substitute it and you get (-2 -3)(-2 +2) = -5 * 0 = 0

OpenStudy (deadshot):

Oh, ok, now I get it, Thanks!

OpenStudy (anonymous):

so whats the answer because 0 is not one of the answer choices